OpenStudy (anonymous):

The solution of a certain differential equation is of the form y(t) = a e^(2 t) + b e^(5 t), where a and b are constants. The solution has initial conditions y(0) = 3 and y'(0) = 5 . Find the solution by using the initial conditions to get linear equations for a and b. y(t)= ?

OpenStudy (turingtest):

you never replied to my earlier efforts to help you... where are you stuck?

OpenStudy (anonymous):

I thought I had it figured out but the answer didn't match

OpenStudy (anonymous):

I don't know where I went wrong

OpenStudy (turingtest):

\[y(t)=Ae^{2t}+Be^{5t}\]and we also need \(y'(t)\) what did you get for that?

OpenStudy (turingtest):

\[y'(t)=2Ae^{2t}+5Be^{5t}\]so what system results from subbing in \(t=0\) ?

OpenStudy (anonymous):

Yup that's what I used, e^0 = 1 so 2A + 5B

OpenStudy (turingtest):

well first off what did you get for a and b ?

OpenStudy (anonymous):

Just A + B right?

OpenStudy (turingtest):

=what?

OpenStudy (anonymous):

For y(t), then y'(t) = 2A +5B

OpenStudy (turingtest):

not quite\[y(t)=Ae^{2t}+Be^{5t}\]it's important to make the distinction between that and the given value\[y(0)=A+B=3\]only by using the second part entirely can we solve the system

OpenStudy (turingtest):

by "the second part" I mean both the x and y values of the initial conditions

OpenStudy (turingtest):

so we also have that\[y'(0)=2A+5B=?\]

OpenStudy (anonymous):

so y(0)=A+B=3 and y'(0)= 2A+5B=5

OpenStudy (turingtest):

good, now you only have to solve the system

OpenStudy (anonymous):

This is where I went wrong

OpenStudy (turingtest):

what were your final answers?

OpenStudy (anonymous):

A=1 B=4 ... I think I'm confusing this

OpenStudy (turingtest):

no that's not right I think you may need a system solving review

OpenStudy (turingtest):

let's use elimination: try subtracting \[y'(0)-2[y(0)]\]

OpenStudy (turingtest):

what is the resulting equation?

OpenStudy (anonymous):

a= 8/3 b= 3-(8/3) does this work?

OpenStudy (turingtest):

no, I think you got B turned negative somehow and messed up A...

OpenStudy (anonymous):

b=1/3

OpenStudy (turingtest):

actually b=-1/3

OpenStudy (turingtest):

I suppose that's where you went wrong

OpenStudy (turingtest):

...are you not seeing it?

OpenStudy (anonymous):

I'm using a=8/3 and b=-1/3 and its still not working

OpenStudy (turingtest):

that's because \[A\neq\frac83\]

OpenStudy (turingtest):

use the first equation and remember the signs that's what's getting you, those darn minus signs

OpenStudy (turingtest):

\[y'(0)=2A+5B=5\]\[-2y(0)=-2A-2B=-6\]so\[y'(0)-2y(0)\implies3B=-1\implies B=-\frac13\]now use the formula for \(y(0)\) again...

OpenStudy (turingtest):

\[y(0)=A+B=3\implies A=3-B\]

OpenStudy (anonymous):

And isn't 3-(1/3) = 8/3

OpenStudy (turingtest):

but\[B\neq\frac13\]watch your signs!!!

OpenStudy (anonymous):

Ohhhh! My bad lol, so it's 10/3

OpenStudy (turingtest):

yeah :) and if these things are allowed you could have just solved the system using wolfram http://www.wolframalpha.com/input/?i=sovle%20x%2By%3D3%2C2x%2B5y%3D5&t=crmtb01 that becomes reasonable with larger systems

OpenStudy (anonymous):

That's where I went wrong :/ Thanks for your patience :-)

OpenStudy (anonymous):

Ahh thanks for the link, it'll be useful!

OpenStudy (turingtest):

...but you should know how to do a 2x2 system like this you're welcome, anytime :)